OpenStudy (anonymous):

If the numerator of a fraction is increased by 3, the fraction becomes 3/4. If the denominator is decreased by 7, the fraction becomes 1. What is the numerator of the fraction? 2 9 16

OpenStudy (anonymous):

?

OpenStudy (anonymous):

your fraction: x/y (x+3)/y=3/4 x/(y-7)=1 4(x+3)=3y x=y-7 4x-3y=-12 x-y=-7 solve this system to find x and y x=y-7 substitutung in 1º equation: 4y-28 -3y=-12 y=16 x=9